OpenStudy (firejay5):

Simplify each expression. Show work and explain. 24. 3 over 5d / (-9) over (15df) 25. p^3 over 2q / (-p) over 4q 26. 2x^3y over z^5 / (4xy over z^3)^2 27. xy over a^3 / (xy over ab)^3

jimthompson5910 (jim_thompson5910):

\[\Large \frac{3}{5d} \div \frac{-9}{15df}\] is the same as \[\Large \frac{3}{5d} \times \frac{15df}{-9}\]

jimthompson5910 (jim_thompson5910):

from there you multiply straight across and you simplify as much as possible

OpenStudy (firejay5):

@phi

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (firejay5):

What about #25

jimthompson5910 (jim_thompson5910):

use the same idea: flip the second fraction and multiply then reduce as much as possible

OpenStudy (firejay5):

I did and I am stuck. I have so far 4p^3q over (-2pq)

jimthompson5910 (jim_thompson5910):

p^3 divided by p = ???

jimthompson5910 (jim_thompson5910):

oh before that 4 divided by -2 = ???

jimthompson5910 (jim_thompson5910):

and finally q/q =??